ON THE ORIENTATION OF SUBSTITUENTS IN N-METHYL- and N, N-DIMETHYL- ANILINES IN THE EXCITED STATES

\underline{NARASIMHA H. AYACHIT}, G. NEERAJA RANI, Department of Physics, SDM College of Engineering and Technology, Dharwad, Karnataka, India.

The orientation of amino group in aniline was considered to change to 180 degrees in the excited state from 38 degrees in the ground state by Suppan[1], to determine the excited electric dipole moment using solvent shifts. Ayachit[2] freed the problem of assumption of considering the orientation in the excited state. This method led to not only to the better calculation of excited state dipole moments also gave a novel method of calculating the orientation of substituents in mono substituted benzenes in their excited states. This method was shown to work successfully in some other cases [3, 4]. In this paper the method has been applied to N-methyl- and N, N-dimethyl- anilines and the orientation of the substituents in the excited state seems to follow the same changes as in amino group in aniline. (References:1. P. Suppan, Chemical Physics Letters 94, 272(1983): 2. N. H. Ayachit, Chemical Physics Letters 164, 272(1989).; 3. N. H. Ayachit, D. K. Deshpande, M. A. Shashidhar & K. Suryanarayana Rao, Spectrochimica Acta, 42A, 585, 1405(1986); 4. N. H. Ayachit & G. Neeraja Rani, Physics and Chemistry of Liquids, 45, 41(2007))